OpenStudy (anonymous):

Sara's gross pay is $600, from which $27 is deduced for OASDI, $8 for Medicare, and $10 for income tax. What is her net pay?

OpenStudy (anonymous):

600 - 27 - 8 - 10 = 573 - 8 - 10 = 565 - 10 = 555

OpenStudy (anonymous):

so why did you do 573-8-10=565